30g of metal at 90oC is placed at 20oC in a calorimeter of mass 70g made of the same metal. If the final temperature is 22.1oC. Calculate the specific heat capacity of the metal.


1
Expert's answer
2022-01-02T18:24:13-0500

"Q_1=Q_2,"

"Q_1=cm_1(t_1-t),"

"Q_2=cm_2(t-t_2),"

"cm_1(t_1-t)=cm_2(t-t_2),"

"c=920~\frac J {kg\cdot °C}."